OpenStudy (anonymous):

between what two consecutive integers does the square root 71 fall between

OpenStudy (anonymous):

heloo someone help

OpenStudy (lukebluefive):

Calculate the square root of 71 using a calculator or something. The answer is the two integers it is between.

OpenStudy (anonymous):

8 and 9

OpenStudy (whpalmer4):

Make a table of the integers and their squares: \(x\) \(x^2\) 0 0 1 1 2 4 3 9 etc. Continue until you've gone past 71 in the right hand column. Then look to see which two numbers on the left gave values in the right column just smaller and larger than 71.

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

what about 87

OpenStudy (jdoe0001):

uze das kalkulattor

OpenStudy (jdoe0001):

\(\bf {\color{red}{ \square }}...\sqrt{87}...{\color{red}{ \square }}\)

OpenStudy (anonymous):

what 2 numbers is it

OpenStudy (whpalmer4):

Use the same procedure I gave you.

OpenStudy (anonymous):

\[8^2=64,9^2=81\]

OpenStudy (anonymous):

whx

OpenStudy (anonymous):

\[\left( \sqrt{71} \right)^2=71\]